Consider two hosts, A and B, are separated by 2,000 kilometers and are directly connected by a link with transmission rate R=15Mbps. The propagation speed over the link is 2.0*108 meters/sec. What is the propagation delay of the link?
Answer #1

Propagation speed = 2*10^8 meters/sec

Therefore, Propagation delay = link length / signal speed

= (2000*10^3)/(2*10^8 )sec = 0.01 sec